# Aesthetic Post-Training Diffusion Models from Generic Preferences with Step-by-step Preference Optimization

### v2

@@ 5.2 Main Evaluation on Aesthetic Alignment
Automatic metrics. Using the four automatic scores (Section 5.1), we compare SPO with Diffusion-DPO, D3PO, and DDPO in Table 1 and Table 2 based on SDXL and SD-1.5, respectively. Note that ImageReward, PickScore, and HPSV2 assess overall image quality including aesthetics.
@@ 5.3 Evaluation on Image-Text Alignment
SPO has mixed results compared with Diffusion-DPO. Table 1 and Table 2 demonstrate improvement over Diffusion-DPO, and yet our user studies support a performance tie or SPO’s slight win in prompt alignment. Further, GenEval results indicate that SPO is not as good as Diffusion-DPO in improving image-text alignment, which we tend to agree on, because the SPO design does not fully capture layout changes. But an interesting insight is that image-text alignment and aesthetics are probably often blended. So while SPO is designed for aesthetic alignment, it still yields some prompt alignment improvement over SDXL, but the improvement is not as much as Diffusion-DPO.
@@ 5.4 Further Analysis
Figure 7: Qualitative comparison between Glyph-SDXL and Glyph-SDXL + SPO in graphic design image generation. SPO consistently improves image aesthetics by creating nuanced textures, nicer colors without sacrificing image content accuracy.
Effectiveness of step-aware preference model (SPM). An important design of SPM is step conditioned noise addition. To verify its usefulness, we remove the time condition in SPM and use uniform noise sampling for all steps (SPM w/o step). A second variant is to simply use the PickScore model [11] which has no time condition and is trained on clean images only. Results in Table 9 show that both variants lead to performance drop, validating the SPM design.
Effectiveness of random selection for initializing next iteration. At the end of each SPO iteration, we need to initialize the next iteration, where random selection is an option. Here we compare random selection with using the win sample $\boldsymbol{x}^{w}_{t-1}$ or the lose sample $\boldsymbol{x}^{l}_{t-1}$ for initialization. Results are presented in Table 9. We clearly find that random selection is better. If we only selecting $\boldsymbol{x}^{w}_{t-1}$ or $\boldsymbol{x}^{l}_{t-1}$, it causes training to be biased towards the intermediate samples that are more preferred or more dispreferred, respectively. This prevents the network from learning from more diverse trajectories, deteriorating generalization ability.
@@
First, when increasing $k$, discrepancy of sampled images becomes larger (but still small enough to only reflect image detail differences). The larger contrast between the preferred and dispreferred samples benefits the model in learning human preference. Second, when $k$ is too large, quality of the most dispreferred sample would be lower than average of samples generated by the base model. As a result, the “push away” effect of the dispreferred image is weakened, causing performance degradation. We choose $k=4$.
@@
Computational cost. We use 4×4$\times$4 × A100 GPUs, which take $12$ and $29.5$ hours to fine-tune SD-1.5 and SDXL, respectively. We also spend $8$ and $29$ hours training SPM for SD-1.5 and SDXL, respectively. In comparison, the GPU hours used for fine-tune SD-1.5 and SDXL using Diffusion-DPO are $384$ and $4,800$, respectively. As a result, the total training GPU hours of SPO-SD-1.5 and SPO-SDXL is about 20.8% and 4.9% of DPO-SD-1.5 and DPO-SDXL, respectively. This significant efficiency gain is probably because of the SPO design, where the image details are properly highlighted for the network to learn.
Generalization to text generation. We verify the generalization of SPO by simply marrying the LoRA weights of SPO-SDXL to the Glyph-SDXL model [14], which specializes in design image generation. Qualitative examples are shown in Fig. 7, where we observe that SPO consistently improves the visual appeal of Glyph-SDXL images, e.g., richer texture of the elephant, flower, robot, and mug.
@@ 6 Conclusion
This paper studies how to align diffusion models with human aesthetic preference. This problem is challenging for two reasons. First, existing two-trajectory methods exhibit large image discrepancies, preventing models from focusing on aesthetic nuances. Second, it is very non-trivial to collect aesthetic-only preference data, while existing datasets record general preferences that may conflict with image aesthetic preferences. To improve aesthetic alignment, we propose an aesthetic alignment solution that distills aesthetics from generic preference data. Specifically, the proposed step-by-step preference optimization method allows a pair of samples to originate from the same image, so their differences are relatively small after one or very few steps, which would reflect their image details or aesthetics. Our preference model captures these differences, enabling the model to improve towards generating better image details. In experiments, we demonstrate that SPO better aligns SD-1.5 and SDXL with human aesthetic preference compared with DPO-based methods and is efficient to train.
@@
Figure 8: Image samples showing disagreement between general preference and aesthetic preference. These images are generated by SDXL. The win trajectories in both examples have inferior aesthetics / details, which are detailed in Section 1.
Figure 9: More image samples showing disagreements between general preference and aesthetic preference in the Pick-a-Pic dataset (best viewed when zoomed in). These images were generated by various diffusion models. Prompt of (a): “a stuffed animal of a blue fox”. Prompt of (b): “girl wearing red lipstick and black leggings”. Prompt of (c): “4 cars racing”. Prompt of (d): “gangsta clothed chicken”. Prompt of (e): “Equity markets were mixed Monday”. Prompt of (f): “A Kludde. A mythical monstrous black furry nocturnal dog with bear claws, green glistening scaled wings and glowing crimson eyes. Several heavy chains hang from its body and ankles”. Prompt of (g): “**a portrait of a 3D cockroach, wearing a bitcoin shirt, in Hawaii, on the beach, hyper-realistic, ultra-detailed, photography, hyper-realistic, photo-realistic, ultra-photo-realistic, super-detailed, intricate details, 8K, surround lighting, HDR”. Prompt of (h): “a suit of armour constructed from meat”.
@@
1 More Examples of Disagreement Between General Preferences and Aesthetic Preferences
@@
We also examine the preference pairs from the Pick-a-Pic V1 dataset and showcase examples of disagreements between the Pick-a-Pic general preference labels and aesthetic preference labels in Figure 9. These disagreements in the dataset hinder model’s improvement in aesthetics.
@@ 2 Timestep-conditional CLIP Vision Encoder
We initialize the step-aware preference model (SPM) with PickScore [11] weights. To preserve the pretrained knowledge, the weights of the linear layer responsible for generating $\gamma$, $\beta$, and $\alpha$ are initialized such that $\gamma=0$, $\beta=0$, and $\alpha=1$, ensuring that the SPM’s output matches the pretrained PickScore model’s output at the beginning of training.

### v3

@@ 5.2 Main Evaluation on Aesthetic Alignment
Automatic metrics. Using the four automatic scores (Section 5.1), we compare SPO with Diffusion-DPO, D3PO, etc., in Table 1 and Table 2 based on SDXL and SD-1.5, respectively. Note that ImageReward, PickScore, and HPSV2 assess overall image quality including aesthetics.
@@ 5.3 Evaluation on Image-Text Alignment
SPO has mixed results compared with Diffusion-DPO. Table 1 and Table 2 demonstrate improvement over Diffusion-DPO, and yet our user studies support a performance tie or SPO’s slight win in prompt alignment. Further, GenEval results indicate that SPO is not as good as Diffusion-DPO in improving image-text alignment, which we tend to agree with, because the SPO design does not fully capture layout changes. But an interesting insight is that image-text alignment and aesthetics are probably often blended. So while SPO is designed for aesthetic alignment, it still yields some prompt alignment improvement, but the improvement is not as much as Diffusion-DPO.
@@ 5.4 Further Analysis
Figure 7: Qualitative comparison between Glyph-ByT5-SDXL and Glyph-ByT5-SDXL + SPO in graphic design image generation. SPO consistently improves image aesthetics by creating nuanced textures and vibrant colors without sacrificing image content accuracy.
Effectiveness of step-aware preference model (SPM). An important design of SPM is the addition of timestep conditioning. To verify its usefulness, we remove the timestep conditioning in SPM (SPM w/o step). A second variant is to simply use the PickScore model [11] which has no time condition and is trained on clean images only. Results in Table 9 show that both variants lead to performance drop, validating the SPM design.
Effectiveness of random selection for initializing next iteration. At the end of each SPO iteration, we need to initialize the next iteration, where random selection is an option. Here we compare random selection with using the win sample $\boldsymbol{x}^{w}_{t-1}$ or the lose sample $\boldsymbol{x}^{l}_{t-1}$ for initialization. Results are presented in Table 9. We clearly find that random selection is better. If we only select $\boldsymbol{x}^{w}_{t-1}$ or $\boldsymbol{x}^{l}_{t-1}$, training becomes biased towards the intermediate samples that are more preferred or more dispreferred, respectively. This prevents the network from learning from more diverse trajectories, deteriorating generalization ability.
@@
First, when increasing $k$, the discrepancy within sampled pairs becomes larger (but still small enough to only reflect image detail differences). The larger contrast between the preferred and dispreferred samples helps the model learn human preferences. Second, when $k$ is too large, quality of the most dispreferred sample would be lower than average of samples generated by the base model. As a result, the “push away” effect of the dispreferred image is weakened, causing performance degradation. We choose $k=4$.
@@
Computational cost. We use 4×4$\times$4 × A100 GPUs, which take $12$ and $29.5$ hours to fine-tune SD-1.5 and SDXL, respectively. We also spend $8$ and $29$ hours training SPM for SD-1.5 and SDXL, respectively. In comparison, the GPU hours used for fine-tuning SD-1.5 and SDXL using Diffusion-DPO are $384$ and $4,800$, respectively. As a result, the total training GPU hours of SPO-SD-1.5 and SPO-SDXL is about 20.8% and 4.9% of DPO-SD-1.5 and DPO-SDXL, respectively. This significant efficiency gain is probably because of the SPO design, where the image details are properly highlighted for the network to learn.
Generalization to Text Generation. We verify the generalization of SPO by simply marrying the LoRA weights of SPO-SDXL to the Glyph-ByT5-SDXL model [14], which specializes in design image generation. Qualitative examples are shown in Fig. 7, where we observe that SPO consistently improves the visual appeal of Glyph-ByT5-SDXL images, e.g., richer texture of the elephant, flower, robot, and beer mug, while preserving the text generation ability of Glyph-ByT5-SDXL.
@@ 6 Conclusion
This paper studies how to align diffusion models with human aesthetic preference. This problem is challenging for two reasons. First, existing two-trajectory methods exhibit large image discrepancies, preventing models from focusing on aesthetic nuances. Second, it is very non-trivial to collect aesthetic-only preference data, while existing datasets record generic preferences that may conflict with image aesthetic preference. To improve aesthetic alignment, we propose an aesthetic alignment solution that distills aesthetics from generic preference data. Specifically, the proposed step-by-step preference optimization method allows a pair of samples to originate from the same image, so their differences are relatively small after one or very few steps, which would reflect their image details or aesthetics. Our preference model captures these differences, enabling the model to improve towards generating better image details. In experiments, we demonstrate that SPO better aligns SD-1.5 and SDXL with human aesthetic preference compared with other DPO methods and is efficient to train.
Acknowledgments
We gratefully acknowledge the support of the ARC Future Fellowship (FT240100820), awarded to Liang Zheng.
@@ Acknowledgments
Figure 8: Image samples showing disagreement between generic preferences and aesthetic preference. These images are generated by SDXL. The win trajectories in both examples have inferior aesthetics / details, which are detailed in Section 1 of the main text.
Figure 9: More image samples showing disagreements between generic preferences and aesthetic preference in the Pick-a-Pic dataset (best viewed when zoomed in). These images were generated by various diffusion models. Prompt of (a): “a stuffed animal of a blue fox”. Prompt of (b): “girl wearing red lipstick and black leggings”. Prompt of (c): “4 cars racing”. Prompt of (d): “gangsta clothed chicken”. Prompt of (e): “Equity markets were mixed Monday”. Prompt of (f): “A Kludde. A mythical monstrous black furry nocturnal dog with bear claws, green glistening scaled wings and glowing crimson eyes. Several heavy chains hang from its body and ankles”. Prompt of (g): “**a portrait of a 3D cockroach, wearing a bitcoin shirt, in Hawaii, on the beach, hyper-realistic, ultra-detailed, photography, hyper-realistic, photo-realistic, ultra-photo-realistic, super-detailed, intricate details, 8K, surround lighting, HDR”. Prompt of (h): “a suit of armour constructed from meat”.
@@
1 More Examples of Disagreements Between Generic and Aesthetic Preferences
@@
We also examine the preference pairs from the Pick-a-Pic V1 dataset and showcase examples of disagreements between the Pick-a-Pic generic preference labels and aesthetic preference labels in Figure 9. These disagreements in the dataset hinder model’s improvement in aesthetics.
@@ 2 Timestep-conditional CLIP Vision Encoder
We initialize the step-aware preference model (SPM) with PickScore [11] weights. To preserve the pretrained knowledge, the weights of linear layers responsible for generating $\gamma$, $\beta$, and $\alpha$ are initialized such that $\gamma=0$, $\beta=0$, and $\alpha=1$, ensuring that the SPM’s output matches the pretrained PickScore model’s output at the start of training.